OpenStudy (anonymous):

find the limit lim cos 4x / x -> - infinity x

OpenStudy (anonymous):

cos4x goses to 1 so u have 1/infinity = infinity?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty }\frac{ \cos4x }{ x } \]

sam (.sam.):

The x in cosx will just alternate -1 to 1, so the only one that grows is the denominator, so you get zero for that

OpenStudy (anonymous):

yea. my mistake

OpenStudy (anonymous):

okay. so x is just = 0.

sam (.sam.):

\[\lim_{x \rightarrow \infty}\frac{\cos(4x)}{x}=0\]